OpenStudy (anonymous):

How to solve this problem and graph -2×+y=3

OpenStudy (pottersheep):

write it out in y = mx + b form :)

OpenStudy (pottersheep):

so y = 2x + 3

OpenStudy (aum):

-2x + y = 3 add 2x y = 2x + 3 The y-intercept is 3. Find another point by putting x = 1: y = 2(1) + 3 = 5 So (1,5) is another point. The first point is (0, 3). Plot the two points and draw a line using a ruler to pass through those two points.